OpenStudy (cherryfrost27):

I WILL GIVE YOU A MEDAL! PLEASE HELP ME.. The amount of turkey to ham at the deli in the supermarket was 7:3 in the morning. By the end of the day , 15lb. of tukeys had been sold.The ratio of turkey to ham was now 3:2. How many pounds of turkey did the the deli have in the morning?

OpenStudy (cherryfrost27):

@Benji

OpenStudy (nguyenvui):

Suppose that supermarket has not sell any ham yet. In the morning, x = the amount of turkey y = the amount of ham rate of these food \[\frac{ x }{ 7 } = \frac{ y }{ 3 }\] By the end of the day, x - 15 = turkey remaining y = amount of ham (was not sold any) rate of those food \[\frac{ x-15 }{ 3 } = \frac{ y }{ 2 }\] Now u can find the answer Feel free to help!! ^^

OpenStudy (loveneenah):

The answer is y/2+5